A farmer has 50 feet of fencing to enclose a rectangular garden with an area of exactly 144 square feet. The area of the garden can be modeled by the equation x(25-x)-144=0, where x represents the length of the garden. what is the length of the garden?
Answer by CharlesG2(834)

A farmer has 50 feet of fencing to enclose a rectangular garden with an area of exactly 144 square feet. The area of the garden can be modeled by the equation x(25-x)-144=0, where x represents the length of the garden. what is the length of the garden?

x(25 - x) - 144 = 0
25x - x^2 - 144 = 0
x^2 - 25x + 144 = 0
(x - 9)(x - 16) with FOIL = x^2 - 16x - 9x + 144 = x^2 - 25x + 144
x = 9 or x = 16
9 + 9 = 18
16 + 16 = 32
18 + 32 = 50
length is 16 feet, width is 9 feet